# Improving Complex Reasoning over Knowledge Graph with Logic-Aware Curriculum Tuning

### v1

@@ 4.3 Fine-tuning Enhanced by Curriculum Learning
Method $avg_{p}$ $avg_{ood}$ $avg_{n}$ 1p 2p 3p 2i 3i pi ip 2u up 2in 3in inp pin pni FB15K GQE 28.0 20.1 - 54.6 15.3 10.8 39.7 51.4 27.6 19.1 22.1 11.6 - - - - - Query2Box 38.0 29.3 - 68.0 21.0 14.2 55.1 66.5 39.4 26.1 35.1 16.7 - - - - - BetaE 41.6 34.3 11.8 65.1 25.7 24.7 55.8 66.5 43.9 28.1 40.1 25.2 14.3 14.7 11.5 6.5 12.4 CQD-CO 46.9 35.3 - 89.2 25.3 13.4 74.4 78.3 44.1 33.2 41.8 21.9 - - - - - CQD-Beam 58.2 49.8 - 89.2 54.3 28.6 74.4 78.3 58.2 67.7 42.4 30.9 - - - - - ConE 49.8 43.4 14.8 73.3 33.8 29.2 64.4 73.7 50.9 35.7 55.7 31.4 17.9 18.7 12.5 9.8 15.1 GNN-QE 72.8 68.9 38.6 88.5 69.3 58.7 79.7 83.5 69.9 70.4 74.1 61.0 44.7 41.7 42.0 30.1 34.3 QTO 74.0 71.8 49.2 89.5 67.4 58.8 80.3 83.6 75.2 74.0 76.7 61.3 61.1 61.2 47.6 48.9 27.5 LARK 56.1 43.1 18.4 72.8 50.7 36.2 66.9 60.4 56.1 23.5 52.4 40.6 16.2 5.7 33.7 26.1 10.0 LACT 82.6 71.9 56.9 93.5 73.5 59.6 92.3 82.3 76.8 75.9 74.6 60.4 81.2 61.6 52.0 43.5 41.7 FB15K-237 GQE 16.3 10.3 - 35.0 7.2 5.3 23.3 34.6 16.5 10.7 8.2 5.7 - - - - - Query2Box 20.1 15.7 - 40.6 9.4 6.8 29.5 42.3 21.2 12.6 11.3 7.6 - - - - - BetaE 20.9 14.3 5.5 39.0 10.9 10.0 28.8 42.5 22.4 12.6 12.4 9.7 5.1 7.9 7.4 3.5 3.4 CQD-CO 21.8 15.6 - 46.7 9.5 6.3 31.2 40.6 23.6 16.0 14.5 8.2 - - - - - CQD-Beam 22.3 15.7 - 46.7 11.6 8.0 31.2 40.6 21.2 18.7 14.6 8.4 - - - - - FuzzQE 24.0 17.4 7.8 42.8 12.9 10.3 33.3 46.9 26.9 17.8 14.6 10.3 8.5 11.6 7.8 5.2 5.8 ConE 23.4 16.2 5.9 41.8 12.8 11.0 32.6 47.3 25.5 14.0 14.5 10.8 5.4 8.6 7.8 4.0 3.6 GNN-QE 26.8 19.9 10.2 42.8 14.7 11.8 38.3 54.1 31.1 18.9 16.2 13.4 10.0 16.8 9.3 7.2 7.8 QTO 33.5 27.6 15.5 49.0 21.4 21.2 43.1 56.8 38.1 28.0 22.7 21.4 16.8 26.7 15.1 13.6 5.4 LARK 50.7 41.0 10.6 73.6 40.5 26.8 46.1 43.1 49.9 22.9 62.8 28.3 6.5 3.4 23.2 16.5 3.2 LACT 57.0 44.4 21.9 76.5 54.3 30.3 56.0 54.5 54.6 36.9 56.5 29.7 17.6 33.1 27.1 19.8 11.2 NELL995 GQE 18.6 12.5 - 32.8 11.9 9.6 27.5 35.2 18.4 14.4 8.5 8.8 - - - - - Query2Box 22.9 15.2 - 42.2 14.0 11.2 33.3 44.5 22.4 16.8 11.3 10.3 - - - - - BetaE 24.6 14.8 5.9 53.0 13.0 11.4 37.6 47.5 24.1 14.3 12.2 8.5 5.1 7.8 10.0 3.1 3.5 CQD-CO 28.8 20.7 - 60.4 17.8 12.7 39.3 46.6 30.1 22.0 17.3 13.2 - - - - - CQD-Beam 28.6 19.8 - 60.4 20.6 11.6 39.3 46.6 25.4 23.9 17.5 12.2 - - - - - FuzzQE 27.0 18.4 7.8 47.4 17.2 14.6 39.5 49.2 26.2 20.6 15.3 12.6 7.8 9.8 11.1 4.9 5.5 ConE 27.2 17.6 6.4 53.1 16.1 13.9 40.0 50.8 26.3 17.5 15.3 11.3 5.7 8.1 10.8 3.5 3.9 GNN-QE 28.9 19.6 9.7 53.3 18.9 14.9 42.4 52.5 30.8 18.9 15.9 12.6 9.9 14.6 11.4 6.3 6.3 QTO 32.9 24.0 12.9 60.7 24.1 21.6 42.5 50.6 31.3 26.5 20.4 17.9 13.8 17.9 16.9 9.9 5.9 LARK 52.9 26.9 12.4 87.8 45.7 33.5 51.3 48.7 23.1 22.2 20.6 41.1 9.9 5.9 24.5 13.3 7.3 LACT 60.1 32.0 17.2 91.4 53.6 40.6 62.2 54.9 31.4 34.8 27.0 34.0 16.0 21.2 21.0 16.3 11.6
Table 1: Test MRR results (%) on complex query answering across all query types. $avg_{p}$ is the average on EPFO queries; $avg_{ood}$ is the average on out-of-distribution (OOD) queries, which consist pi/ip/2u/ip queries; $avg_{n}$ is the average on queries with negation.
4.4 Reasoning Module
@@
We use the final LACT LLM as the answer generator, as shown in Figure 1(c). We retrieve relevant information and textualize the FOL query, and finally we populate it into the template in 4.4 to generate responses.
@@
We use the LLM to do a simple text generation task to get the answer. After fine-tune, LACT LLM can follow the output mode in training stage in Figure 1, so we can extract final answers through simple regular expressions with the template in 4.4.
@@
5 Experiments
5.1 Training Settings
Training datasets We opt for the the most popular datasets: FB15K, FB15K-237, NELL995. Detailed information about dataset is listed in  D.1. We used the training set of the above dataset as the original training data.
Training Details We use open-source model LLaMA-2-base, including two different parameter sizes: 7B and 13B, as the base model for fine-tuning. All LLaMA-2-7B and LLaMA-2-13B models are trained by fully fine-tuning.
@@ 5.1 Training Settings
Training Cost We trained for a total of approximately 10 hours using 8*A100 during three training stages.
5.2 Experimental Settings
Baseline Methods For comparing with KGE, we chose the following representative methods as baselines: GQE (?), Query2Box(Q2B) (?), BetaE (?), CQD (?), ConE (?), GNN-QE (?), QTO (?). We also compared our method to an LLM-based method, LARK  (?).
Evaluation Protocol Following previous works (?), we use mean reciprocal rank (MRR) as standard evaluation protocols for evaluating complex reasoning over knowledge graph. In the filtered setting, all easy and hard answers are filtered out during ranking. The detail could be found in Appendix  D.3.
5.3 Main Results
The main experiment results of three datasets are shown in Table 1. The baselines were trained on 1p/2p/3p/2i/3i queries, hence other than these 4 types of EPFO queries serve as OOD queries, and we report the average result on these queries in $avg_{ood}$. We observe that LACT significantly outperforms baseline methods across all datasets. Notably, LACT yields an averag gain of 7.3%, 2.9%, and 6.3% on $avg_{p}$, $avg_{odd}$, and $avg_{n}$, compared to the previous SOTA method, especially more challenging datasets like FB15K-237 and NELL995. This suggests that our method has better reasoning capability and captures a broad range of relations to effectively utilize this capability for enhancing the performance of complex queries.
Table 2: Accuracy results (%) of whether to use BTD on hard complex query answering across all query types, evaluated on FB15k.
Table 3: Accuracy results (%) of whether to use CL on hard complex query answering across all query types, evaluated on FB15k.
@@ 5.3 Main Results
Table 4: MRR results (%) based on differnet base models across all query types, evaluated on FB15k-237.
5.4 Ablation Studies
To verify the effectiveness of the LACT design, we conduct a two-part ablation study. The first part is designed to verify the effectiveness of logical chain decomposition and the second part is designed to verify the effectiveness of curriculum learning.
From a training perspective, as shown in Figure 2, although perplexity (PPL) and training loss of decomposed queries before training was slightly higher than that of ordinary queries, we found that as training progresses, the loss and PPL of decomposed queries will quickly decrease to levels much lower than ordinary queries, proving that chain decomposition is effective to reduce the difficulty of learning complex queries.
5.5 Transferability Study
Task-level transferability. The results in Table 1 show that our method achieves a relative gain of 9.9% on the OOD task, which demonstrates the strong generalization of our fine-tuning framework. Even in the OOD queries, as shown in Table 5, more than 95% of test samples can still follow logical chain reasoning. These phenomena indicate strong generalization ability of LACT.
Metric pi ip 2u up Prodecomposed 98.7 100.0 97.8 100.0 Protrue,decomposed 98.6 99.9 97.8 99.6

### v4

@@ Introduction
Method $avg_{p}$ $avg_{ood}$ $avg_{n}$ 1p 2p 3p 2i 3i pi ip 2u up 2in 3in inp pin pni Dataset FB15K GQE 28.0 20.1 - 54.6 15.3 10.8 39.7 51.4 27.6 19.1 22.1 11.6 - - - - - Query2Box 38.0 29.3 - 68.0 21.0 14.2 55.1 66.5 39.4 26.1 35.1 16.7 - - - - - BetaE 41.6 34.3 11.8 65.1 25.7 24.7 55.8 66.5 43.9 28.1 40.1 25.2 14.3 14.7 11.5 6.5 12.4 CQD-CO 46.9 35.3 - 89.2 25.3 13.4 74.4 78.3 44.1 33.2 41.8 21.9 - - - - - CQD-Beam 58.2 49.8 - 89.2 54.3 28.6 74.4 78.3 58.2 67.7 42.4 30.9 - - - - - ConE 49.8 43.4 14.8 73.3 33.8 29.2 64.4 73.7 50.9 35.7 55.7 31.4 17.9 18.7 12.5 9.8 15.1 GNN-QE 72.8 68.9 38.6 88.5 69.3 58.7 79.7 83.5 69.9 70.4 74.1 61.0 44.7 41.7 42.0 30.1 34.3 QTO 74.0 71.8 49.2 89.5 67.4 58.8 80.3 83.6 75.2 74.0 76.7 61.3 61.1 61.2 47.6 48.9 27.5 LARK 56.1 43.1 18.4 72.8 50.7 36.2 66.9 60.4 56.1 23.5 52.4 40.6 16.2 5.7 33.7 26.1 10.0 LACT 82.6 71.9 56.9 93.5 73.5 59.6 92.3 82.3 76.8 75.9 74.6 60.4 81.2 61.6 52.0 43.5 41.7 Dataset FB15K-237 GQE 16.3 10.3 - 35.0 7.2 5.3 23.3 34.6 16.5 10.7 8.2 5.7 - - - - - Query2Box 20.1 15.7 - 40.6 9.4 6.8 29.5 42.3 21.2 12.6 11.3 7.6 - - - - - BetaE 20.9 14.3 5.5 39.0 10.9 10.0 28.8 42.5 22.4 12.6 12.4 9.7 5.1 7.9 7.4 3.5 3.4 CQD-CO 21.8 15.6 - 46.7 9.5 6.3 31.2 40.6 23.6 16.0 14.5 8.2 - - - - - CQD-Beam 22.3 15.7 - 46.7 11.6 8.0 31.2 40.6 21.2 18.7 14.6 8.4 - - - - - FuzzQE 24.0 17.4 7.8 42.8 12.9 10.3 33.3 46.9 26.9 17.8 14.6 10.3 8.5 11.6 7.8 5.2 5.8 ConE 23.4 16.2 5.9 41.8 12.8 11.0 32.6 47.3 25.5 14.0 14.5 10.8 5.4 8.6 7.8 4.0 3.6 GNN-QE 26.8 19.9 10.2 42.8 14.7 11.8 38.3 54.1 31.1 18.9 16.2 13.4 10.0 16.8 9.3 7.2 7.8 QTO 33.5 27.6 15.5 49.0 21.4 21.2 43.1 56.8 38.1 28.0 22.7 21.4 16.8 26.7 15.1 13.6 5.4 LARK 50.7 41.0 10.6 73.6 40.5 26.8 46.1 43.1 49.9 22.9 62.8 28.3 6.5 3.4 23.2 16.5 3.2 LACT 57.0 44.4 21.9 76.5 54.3 30.3 56.0 54.5 54.6 36.9 56.5 29.7 17.6 33.1 27.1 19.8 11.2 Dataset NELL995 GQE 18.6 12.5 - 32.8 11.9 9.6 27.5 35.2 18.4 14.4 8.5 8.8 - - - - - Query2Box 22.9 15.2 - 42.2 14.0 11.2 33.3 44.5 22.4 16.8 11.3 10.3 - - - - - BetaE 24.6 14.8 5.9 53.0 13.0 11.4 37.6 47.5 24.1 14.3 12.2 8.5 5.1 7.8 10.0 3.1 3.5 CQD-CO 28.8 20.7 - 60.4 17.8 12.7 39.3 46.6 30.1 22.0 17.3 13.2 - - - - - CQD-Beam 28.6 19.8 - 60.4 20.6 11.6 39.3 46.6 25.4 23.9 17.5 12.2 - - - - - FuzzQE 27.0 18.4 7.8 47.4 17.2 14.6 39.5 49.2 26.2 20.6 15.3 12.6 7.8 9.8 11.1 4.9 5.5 ConE 27.2 17.6 6.4 53.1 16.1 13.9 40.0 50.8 26.3 17.5 15.3 11.3 5.7 8.1 10.8 3.5 3.9 GNN-QE 28.9 19.6 9.7 53.3 18.9 14.9 42.4 52.5 30.8 18.9 15.9 12.6 9.9 14.6 11.4 6.3 6.3 QTO 32.9 24.0 12.9 60.7 24.1 21.6 42.5 50.6 31.3 26.5 20.4 17.9 13.8 17.9 16.9 9.9 5.9 LARK 52.9 26.9 12.4 87.8 45.7 33.5 51.3 48.7 23.1 22.2 20.6 41.1 9.9 5.9 24.5 13.3 7.3 LACT 60.1 32.0 17.2 91.4 53.6 40.6 62.2 54.9 31.4 34.8 27.0 34.0 16.0 21.2 21.0 16.3 11.6
Table 1: MRR scores (%) on complex reasoning for 14 types of queries. $avg_{p}$ represents the mean score of nine ordinary queries including 1p/2p/3p/2i/3i and pi/ip/2u/ip; $avg_{ood}$ is the mean score of out of distribution (OOD) queries, which consist pi/ip/2u/ip queries; $avg_{n}$ is the mean score of queries including negation operation.
Reasoning Module
@@
We retrieve relevant information and textualize the $EFO_{1}$ query, and finally, we populate it into the template in the Appendix to generate queries.
@@
As shown in Appendix, we use the LLM to do a simple text generation task to get the answer. After fine-tuning, LACT can follow the output mode in the training stage in Figure 1, so we can extract final answers through simple regular expressions with the template in the Appendix. The specific prompt can be found in the Appendix.
Experiments
Training Settings
We opt for the most popular datasets: FB15K, FB15K-237, NELL995. We used the training set of the above dataset as the original training data. Detailed information about the dataset and training set is listed in the Appendix.
Experimental Settings
Baseline Methods For comparing with KGE, we chose the following representative methods as baselines: GQE (Hamilton et al. 2018), Query2Box(Q2B) (Choudhary and Reddy 2023), BetaE (Ren and Leskovec 2020), CQD (Arakelyan et al. 2020), ConE (Zhang et al. 2021), GNN-QE (Zhu et al. 2022), QTO (Bai et al. 2023). Moreover, we also compared our LACT with the LLM-based method LARK  (Choudhary and Reddy 2023).
Table 2: Accuracy results (%) evaluated on FB15k of whether to use BTD for hard complex query on all query types.
Evaluation Protocol Following previous works (Ren, Hu, and Leskovec 2019), we choose the Mean Reciprocal Rank (MRR) with standard evaluation protocol for evaluating complex reasoning on KG. The detailed setting can be found in the Appendix.
Main Results
The main experiment results of three datasets are shown in Table 1. The baseline training datasets contain five different queries including 1p/2p/3p/2i/3i, so another four queries are considered as OOD types, so the mean value of OOD types of queries is recorded as $avg_{ood}$. In our experiments, LACT consistently outperforms baseline methods across all datasets. Notably, LACT yields an average gain of 7.3%, 2.9%, and 6.3% on $avg_{p}$, $avg_{odd}$, and $avg_{n}$, compared to the previous SOTA method, especially more challenging datasets like FB15K-237 and NELL995. Our method exhibits superior reasoning capability and effectively captures a wide range of internal relations, leading to enhanced performance on complex queries.
Table 3: Accuracy results (%) evaluated on FB15k of whether to use CL for hard complex query on all query types.
Ablation Studies
To validate the efficacy of the LACT module, we conduct a two-part ablation study. Firstly, we will investigate the impact of logical chain decomposition, while the second part assesses the effectiveness of curriculum learning.”
From a training perspective, as shown in Figure 3, although perplexity (PPL) and training loss of decomposed queries before training were slightly higher than that of ordinary queries, we found that as training progresses, the loss and PPL of decomposed queries will quickly decrease to levels much lower than ordinary queries, proving that chain decomposition is effective to reduce the difficulty of learning complex queries.
Transferability Study
Task-level transferability. The results in Table 1 show that our LACT achieves a relative gain of 9.9% on the OOD task, which demonstrates the strong generalization of our fine-tuning framework. Even in the OOD queries, as shown in the Appendix, more than 95% of test samples can still follow logical chain reasoning. These phenomena indicate the strong generalization ability of LACT.
Dataset-level transferability. In fact, almost all KGE methods, even if some of the optimization methods claim not to require training, require a KGE base model adapted to a specific dataset, which leads to the inherent defect of extremely poor Transferability of the KGE method. However, as previous research has shown, fine-tuning LLMs is mainly to stimulate the knowledge and capabilities of potentially relevant downstream tasks contained in LLM pre-training. This has also become the theoretical basis for the transferability of fine-tuning methods for LLMs. The results in the Appendix show that the reasoning ability stimulated by one dataset can still be demonstrated in another dataset, which reflects well in the query performance which only dropped less than 5%.
Model-level transferability. We tried analytical experiments with different base models to determine whether our LACT was universal. Obviously, The results in the Appendix show that all types of queries have been improved to a certain extent due to the progress of the base model. Experimental results show that our LACT is suitable for different pedestal models and has strong generalization.
Scalability Study
To verify the scalability of LACT, we scale LACT to different model sizes and data volumes.
Figure 4: (left) MRR performance of LACT and previous SOTA methods at different difficulties based on FB15K. (right) The correlation between related information completeness and accuracy evaluated on FB15K-237, we selected a 3p query and pi query with the same inference path length as the task types. We assume that the completeness of all simple queries is 1.
Discussion
When and Where Does LACT Work?
LACT performs consistently better with more complete information. We take the form of a posteriori that set the completeness of relevant triplets to the proportion of triplets in the inference path of complex reasoning in the provided context, (For example, if the inference path of a 2i query is (Turing Award, winner, Yoshua Bengio), (Canada, citizen, Yoshua Bengio), and we can only retrieve (Turing Award, winner, Yoshua Bengio) in incomplete KG, then our completeness of relevant triplets to 1/2.) and set the completeness of simple queries that can be directly inferred to 1, to obtain the relation between Accuracy and correlation information completeness. As seen in Figure  4, LACT obtains a significant gain when the completeness of relevant information increases, though, with zero relevant information, it remains a certain amount of complex reasoning ability.
LACT performs consistently better on higher difficulties. As mentioned before, we simply divide the difficulty of the query by the number of hops in the query. The results in Figure  4 show that our model yields more gain in tasks of higher-level difficulty and complexity, which benefits from our unique and sophisticated fine-tuning framework.
Conclusion
In this paper, we present a simple and effective fine-tuning framework LACT to boost the complex logical reasoning ability over KGs. LACT is a two-part method that consists of both Binary Tree Decomposition and Curriculum Learning and can be applied to various size LLMs with different data sizes. We empirically demonstrate the effectiveness and universality of the LACT on a series of widely used knowledge graph datasets. Further analyses delve into the underlying mechanism of our LACT, and investigate When and Why Does LACT Work. We hope that our work can inspire more research on combining knowledge graphs and LLMs.
Acknowledgments
This research is supported by NSFC (62225113, 62206202), National Key RD Program of China (2023YFC2705700), The Innovative Research Group Project of Hubei Province under Grants (2024AFA017), the National Research Foundation, Singapore, and the CyberSG R&D Programme Office (“CRPO”), under the National Cybersecurity R&D Programme (“NCRP”), RIE2025 NCRP Funding Initiative (Award CRPO-GC1-NTU-002).
Appendix
Appendix A Neighborhood Retrieval Algorithm
Retrieval Algorithm
For another query, as shown in Figure 6 for each leaf node in DAG, we do depth traversal on the graph. For each step in the traversal process, if this step is a projection, we search for all the possible triplets. Otherwise, we perform corresponding operations on intersection and union respectively to filter out the corresponding entities.
Over-limit Solutions and Token distribution
In all experimental setups, we used the Max Seq Length of 4096, and we simply discarded all out-of-bounds samples and recorded them as 0 at the time of evaluation. In fact, after using our information retrieval algorithm, most of our samples were controlled below 4096. As shown in Figure 7, compared with the algorithm of searching all possible related triples, our retrieval algorithm greatly reduces the consumption of tokens.
Figure 5: Query structure diagram of 14 types of queries
Figure 6: A case on Neighborhood Retrieval Algorithm
Figure 7: Token numbers’ probability distribution in the dataset FB15K-237 and all-context Retrieval Algorithm.
Table 4: Difficulty of different query types where 1 means easy, 2 means medium, 3 means hard.
Appendix B Universal Procedure for Converting from $EFO_{1}$ to Computational Trees
To transform a $EFO_{1}$ formula into its corresponding computation tree, a two-phase process is employed: dependency graph generation, and union branches duplication.
Dependency Graph Generation. Upon encountering a $EFO_{1}$ expression, our primary procedure entails the allocation of distinct nodes to individual variables, while a separate node is assigned to each specific entity for each single-hop query. It is important to acknowledge that multiple nodes may represent one same intermediate entity, given its occurrence among various single-hop atom. Subsequently, undirected edges are employed to establish connections between nodes in accordance with the defined one-hop atoms. In the case where $e^{i}_{j}=r(h,t)$, we link $h$ and $t$ by establishing the edge recorded as $h_{k}$. Similarly, if $e^{i}_{j}=\lnot r(h,t)$, we link $h$ and $t$ by the edge recorded as $\lnot r_{k}$. The variable $k$ serves as a distinguishing label for edges emanating from distinct atom struture. The formulated multi-graph must conform be in a tree-shape, signifying a DAG. Choosing tail node $v_{?}$ to be the root, We define the orientation of edges to consistently direct from child nodes towards their parent nodes, while carefully accommodating inverse relationships. It is important to note that constant entities nherently serve as terminal nodes in the computation tree, as they are each uniquely associated with a solitary variable node.
Union Branches Duplication Subsequently we address the duplication branches within the computation tree. For each path $\tau$ from every leaf node to root, we search for the first node $v_{i}$ which contains the same relations between $v_{i}$ and its possible child nodes $v_{i}^{{}^{\prime}}$, but in different link structure: $r_{i_{1}},r_{i_{2}},\dots,r_{i_{n}}$. These edges were amalgamate into a converged edge $r_{i_{1},i_{2},\dots i_{n}}$ to dissolve the complex multiple edges structure and then they can be amalgamate abide by the following distributive law:
$(A\land B)\lor(A\land C)\Leftrightarrow A\land(B\lor C)$ (7)
Appendix C Prompt Template
The query and output prompt templates can be found in C and C.
@@ Appendix C Prompt Template
Appendix D Difficulty
We divide the difficulty by the number of decomposed subqueries. Query types and their corresponding difficulties are shown in Table 4.
Appendix E Experiment Details
Dataset Details
$\bullet$ FB15K comes from a subset of Freebase, a large knowledge database initiated by Google. FB15k has a total of 592,213 triplets consisting of 15,000 entities and 1,345 relationships.
$\bullet$ FB15K-237 is the subset of FB15k which contains 310,116 triplets consisting of 14,541 entities and 237 relations. FB15K-237 corrects the inverse relationship leakage problem in the test set to a certain extent by selecting a specific relationship subset in FB15K, thereby more realistically reflecting the effect of the model on complex reasoning on the knowledge graph.
$\bullet$ NELL995 is automatically extracted and generated by Carnegie Mellon University’s Never-Ending Language Learning (NELL) project. It was considered as a comprehensive knowledge graph data set that includes 995 relationships in multiple fields.
Training Details
We use open-source model LLaMA-2-base, including two different parameter sizes: 7B and 13B, as the base model for fine-tuning. All LLaMA-2-7B and LLaMA-2-13B models are trained by fully fine-tuning.
@@ Appendix E Experiment Details
Training Cost
We trained for a total of approximately 10 hours using 8*A100 during three training stages.
Query Structure
Taking into account the fairness of the evaluation, we use the same 14 types of queries as in the previous work (Bai et al. 2023). The query structure of each type is shown in Figure 5.
@@
Table 5: MRR results (%) based on differnet base models evaluated on FB15k-237.
Evaluation Protocol Detail
The answers to queries in complex logical reasoning can be divided into simple and hard types, which are distinguished by whether answers of the query can be easily retrieved directly in the knowledge graph. Specifically, in the valid set, we consider answers that can be directly retrieved from the graph in the training set as simple answers, and those that cannot be directly retrieved as hard answers. And in the test set, we consider answers that can be directly retrieved from the valid graph as simple answers, and those that cannot be directly retrieved as hard answers. Referring to previous work (Ren and Leskovec 2020), we choose the Mean Reciprocal Rank (MRR) as the evaluation index on the filtered setting all answers (including easy and hard) would be filtered out when ranking.
Appendix F More Results
Overall comparison
We provide a multi-dimensional comparison with previous state-of-the-art methods in Figure  8
OOD results
1p 2p 3p 2i 3i pi ip 2u up 2in 3in inp pin pni LARK 72.8 50.7 36.2 66.9 60.4 56.1 23.5 52.4 40.6 16.2 5.7 33.7 26.1 10.0 LACT 97.0 87.2 80.8 95.8 85.3 71.7 77.4 85.7 62.6 88.5 81.4 77.8 68.9 65.7
Table 6: Accuracy results (%) evaluated on FB15k on complete KG for complex query on all query types.
Metric pi ip 2u up nin nipn Prodecomposed 98.7 100.0 97.8 100.0 98.4 97.2 Protrue,decomposed 98.6 99.9 97.8 99.6 96.1 96.9